Assume Organic Ice Cream Company, Inc., bought a new ice cream production kit (pasteurizer/homogenizer, cooler, aging vat, freezer, and filling machine) at the beginning of the year at a cost of $16,800. The estimated useful life was four years, and the residual value was $1,280. Assume that the estimated productive life of the machine was 9,700 hours. Actual annual usage was 3,880 hours in Year 1; 2,910 hours in Year 2; 1,940 hours in Year 3; and 970 hours in Year 4.

Required:

1. Complete a separate depreciation schedule for each of the alternative methods.

a. Straight-line.

b. Units-of-production.

c. Double-declining-balance.Complete this question by entering your answers in the tabs below. Req 1A Reg 1B Req 1C Complete a depreciation schedule usinComplete this question by entering your answers in the tabs below. Req 1A Reg 1B Req 10 Complete a depreciation schedule usinComplete this question by entering your answers in the tabs below. Req 1A Req 1B Req 1C Complete a depreciation schedule usin
Answer #1

a) Straight line depreciation

Year Depreciation expense Accumulated depreciation Net book value
At acquisition 16800
1 (16800-1280/4) = 3880 3880 12920
2 3880 7760 9040
3 3880 11640 5160
4 3880 15520 1280

b) Unit of production depreciation

Depreciation expense per unit = (16800-1280/9700) = 1.60

Year Depreciation expense Accumulated depreciation Net book value
At acquisition 16800
1 3880*1.6 = 6208 6208 10592
2 2910*1.6 = 4656 10864 5936
3 1940*1.6 = 3104 13968 2832
4 1552 15520 1280

c) Double decline

Rate = 100/4*2 = 50%

Year Depreciation expense Accumulated depreciation Net book value
At acquisition 16800
1 16800/2 = 8400 8400 8400
2 8400/2 = 4200 12600 4200
3 4200/2 = 2100 14700 2100
4 820 15520 1280